OpenStudy (anonymous):

ln|1+sinx|+ln|1-sinx|=2ln|cosx|

OpenStudy (anonymous):

\(\large\color{black}{ \rm ln|~a~|~~+~~ln|~b~|~~=~~ln|~a\times b~|}\)

OpenStudy (anonymous):

a = 1+sin(x) b = 1-sin(x) apply this rule:)

OpenStudy (anonymous):

Need more clarification?

OpenStudy (anonymous):

Ha! I got here!

OpenStudy (anonymous):

okay, good:) And how do you feel about my initial post? Do you think you understand the rule, and can apply it to your problem?

OpenStudy (anonymous):

I dont understand... can you FOIL 1-sinx ans 1+sinx?

OpenStudy (anonymous):

yes, you can.

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

Just like (a-b)(a+b)=a^2-b^2 so will (1-sin x)( 1 + sin x) equal what?

OpenStudy (anonymous):

2-sin2x?

OpenStudy (anonymous):

why twos?

OpenStudy (jhannybean):

You meant \(1-\sin^2(x)\) yes?

OpenStudy (anonymous):

*1... 1^2 is 1 :D

OpenStudy (jhannybean):

Good.

OpenStudy (anonymous):

and the 2 inside the sin2x is supposed to be the power?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, now we have a rule. (Pythagorean) \(\large\color{black}{ \sin^2(x)+\cos^2(x)=1}\) at first, can you tell me what will we get, IFF we subtract \(\large\color{black}{ \cos^2(x)}\) from both sides (of the rule).

OpenStudy (anonymous):

1-sin2x=cos2x

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

and the you take the 2 to the outside?

OpenStudy (anonymous):

so you already know that it will turn into cos^2x ! good!

OpenStudy (anonymous):

Yes, the 2 goes on the outside as an exponenet.

OpenStudy (anonymous):

(because cos^2x is same as (cos x)^2 )

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

Anytime... you are a very smart student!

OpenStudy (jhannybean):

Remember to medal the user who helped you the most. :)